# SynCDR : Training Cross Domain Retrieval Models with Synthetic Data

### 31 Dec 2023 v1

@@ Abstract
In cross-domain retrieval, a model is required to identify images from the same semantic category across two visual domains. For instance, given a sketch of an object, a model needs to retrieve a real image of it from an online store’s catalog. A standard approach for such a problem is learning a feature space of images where Euclidean distances reflect similarity. Even without human annotations, which may be expensive to acquire, prior methods function reasonably well using unlabeled images for training. Our problem constraint takes this further to scenarios where the two domains do not necessarily share any common categories in training data. This can occur when the two domains in question come from different versions of some biometric sensor recording identities of different people. We posit a simple solution, which is to generate synthetic data to fill in these missing category examples across domains. This, we do via category preserving translation of images from one visual domain to another. We compare approaches specifically trained for this translation for a pair of domains, as well as those that can use large-scale pre-trained text-to-image diffusion models via prompts, and find that the latter can generate better replacement synthetic data, leading to more accurate cross-domain retrieval models. Code for our work is available at https://github.com/samarth4149/SynCDR
Figure 1: Motivation. Cross Domain Retrieval (CDR) problems show up in many applications, and prior work has developed solutions in different scenarios, including when labeled data is absent [18]. What these solutions rely on however, is the presence of same category data in both domains so similar pairs can be discovered during training. When these are missing, such approaches can fail. Our solution is to make up for these missing examples using synthetic data (generated via label preserving translation).
@@ 2 Related Work
Self-supervised Feature Learning for Cross-Domain Retrieval. While approaches above rely on labels for training, these may be expensive in labor-hours to collect. Kim et al. [18], followed by others [47, 11], demonstrated that category discriminative domain-adaptive features can be learned without the presence of any labels using self-supervised criteria within and across domains. This, however, relies on the presence of same categories in both domains in training data. If unavailable, the cross-domain self-supervision losses can degrade retrieval performance by training the model to match unrelated instances. This is the problem we tackle in this paper, by using synthetically generated examples to make up for the missing categories of data.
@@ 3 Approach
Notation. Let the two domains be A and B. At training time, two sets of images are available (with no accompanying labels), $\mathcal{D}_{A}=\{\boldsymbol{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{D}_{B}=\{\boldsymbol{x}^{B}_{i}\}_{i=1}^{n_{B}}$. The ground truth category labels are unknown but let each image in $\mathcal{D}_{A}$ come from a category label in $Y_{A}$, and each image in $\mathcal{D}_{B}$ come from a category label in $Y_{B}$. It is known that $Y_{A}\cap Y_{B}=\phi$ (This is not a strict requirement for SynCDR to function and we report results where $Y_{A}$ and $Y_{B}$ share some categories in Sec. A.2). The outcome of training is a feature extractor $f:\mathcal{X}\rightarrow\mathbb{S}^{d}$ (where $\mathcal{X}$ is the support of all images, and $\mathbb{S}^{d}$ the $d$-dimensional unit hypersphere $\mathbb{S}^{d}=\{\boldsymbol{x};\boldsymbol{x}\in\mathbb{R}^{d},\left|\left|x\right|\right|_{2}=1\}$), such that cosine similarities in the output feature space reflect category semantics.
For evaluation are provided two sets $\mathcal{D}^{(test)}_{A}$ and $\mathcal{D}^{(test)}_{B}$ both of which have images with known category labels and each set has at least one image from each label in $Y_{A}\cup Y_{B}$. At test time, given an image $\boldsymbol{x}^{A}$ in $\mathcal{D}^{(test)}_{A}$, it is desired to fetch images of the same category label from $\mathcal{D}^{(test)}_{B}$, and vice-versa for a given image in domain B. This is done by ranking in decreasing order $\{f(\boldsymbol{x}^{A})^{\top}f(\boldsymbol{x}^{B});\boldsymbol{x}^{B}\in\mathcal{D}^{(test)}_{B}\}$ and picking the top K to calculate precision@K (for different values of K).
@@
Figure 2: Synthetic examples from different translation methods. We compare synthetic data generated using 4 methods. Since ELITE is not specifically restricted to closely mimicking the original image, we can generate more natural examples of target domain data, which serves as better synthetic replacement for missing real data. For more discussion, refer to 3.1.
@@
$L_{PPP}(X_{A},X_{syn,B})$
$=\frac{1}{m}\sum_{\boldsymbol{x}^{A}\in X_{A}}-\log\left(\frac{\exp(f(\boldsymbol{x}^{A})^{\top}f(\boldsymbol{x}_{syn}^{B}))}{\sum_{\boldsymbol{x}\in X_{syn,B}}\exp(f(\boldsymbol{x}^{A})^{\top}f(\boldsymbol{x}))}\right)$
@@
Training SynCDR. For training we combine $L_{PPP}$ with the CDS criterion [18] ($L_{CDS}$), and minimize the following loss via minibatch stochastic gradient descent. To meet length constraints, we excluded a precise definition of $L_{CDS}$ and refer readers to Sec 3 and Fig 3 of [18]. For the sake of current discussion, we simply describe it as a combination of in-domain and cross-domain contrastive criteria given two sets of unlabeled images from two domains.
@@ 3.1 Methods for Synthetic Data Generation
Img2Img/SDEdit [20] uses a pre-trained Stable Diffusion model for image editing by partially following the diffusion forward noising process (so that not all content in the input image is lost due to noising) and then denoising based on a text prompt. For instance in the example in Fig. 2, for translating real images of birds from the CUB dataset to paintings, we used the prompt “A painting of a bird.” and for translating sketches to paintings for Domainnet objects, we used “A painting of an object.”. The method allows for specification of edit strength $\in[0,1]$, which is the fraction of diffusion noising steps to follow. The value 00 leads to no edit and $1$ leads to completely following the prompt and retaining none of the input image information. In our experiments, we set this parameter to $0.3$ based on validation performance. For prompts used for all datasets and other generation parameters, we refer readers to Appendix B.
@@
InstructPix2Pix [2] is a Stable-Diffusion model fine-tuned and adapted for following natural language instructions for editing images. For converting to paintings (for both photos of birds as well as sketches of objects from Domainnet), we input the original image and the prompt “Convert to a painting.”. With InstructPix2Pix, we can provide a parameter $image\_guidance\_scale\geq 1$. The larger the value of this parameter, more is the output image similar to the input image. In our experiments, we set this parameter to $1.2$ based on validation performance. Full prompts and other generation parameters for all datasets and additional examples are in Appendix B.
@@
The above translation methods attempt to closely resemble the original image in translations and we find that this is often not enough to accurately represent data in the other domain. In such scenarios, images may need additional content to make it look like the target domain. For instance, in Fig. 2, while the above methods do a decent job of converting paintings to pencil sketches or real photos to paintings, the translation in the opposite direction is much poorer. In translating a sketch to a painting (row 2), Img2Img and InstructPix2Pix fail to add color and CUT adds some incorrect colors based on what it could learn from unpaired images. Similar is the case in converting a painting of a bird to a realistic photograph (row 4), where a background may be expected but above methods fail to generate one (CUT does generate one, but does not generate an natural looking photograph). While Img2Img and InstructPix2Pix allow for increasing the amount of edit via a controllable parameter, in the supplement, we show that this often leads to non-preservation of image category. The following generation method can avoid these issues.
@@
ELITE [43] is a personalization approach based on Stable Diffusion and a CLIP-based image encoder. The encoder encodes an object instance from an image as a word in Stable Diffusion’s vocabulary, which can then be used for generation with prompts. In the example in Fig. 2, for generating a painting (from either a real photo or a sketch), we input the image to ELITE’s image encoder and use the prompt “A painting of $\langle S\rangle$.” (where $\langle S\rangle$ is the image encoded as a word, or more precisely, a token embedding). For full details, we refer readers to Appendix B.
@@ 4 Experiments
Table 1: Cross Domain Retrieval Prec@1 in different scenarios of DomainNet. Using simply the In-domain ID criterion for training improves performance over ImageNet pre-training. It is also better than CDS since its cross-domain matching criterion cannot find same class instances in the other domain. We find synthetic data can make up for these missing examples as seen in SynCDR’s performance. Additionally, we find that ELITE leads to best quality synthetic data for training visual representations for cross-domain retrieval.
Datasets and Evaluation. We ran experiments on three different datasets : DomainNet [22], CUB [41, 42] and Office-Home [39]. We experimented with the 126 class subset of DomainNet used in [18] with 3 domains : clipart, painting and sketch. For CUB, we use two domains : Real (coming from the CUB-200-2011 data [41]) and Paintings (coming from [42]). In CUB, we removed categories that had fewer than 5 paintings, leaving us with 187 categories. The Office-Home dataset has 65 categories of objects in 4 domains : Art, Clipart, Product and Real World.
We experimented with each possible pair of domains in each dataset. For the experiments, we made a 50:20:30 split of the data in the two domains using the first for training, second for validation and third for testing. For each pair of domains we split the training data into two disjoint sets of categories (of equal size) and run two experiments : one where category set 1 is from domain A and category set 2 is from domain B, and the other where this is reversed. Note that testing in either of these cases takes place on the same data, and we report average Prec@1 of retrieval in either direction (given domain A image, fetch a domain B image of same category and vice-versa), and the overall average of these over all pairs of domains in the dataset. In Sec. A.4, we also report Prec@5 and Prec@15.
Other Implementation Details. Similar to CDS [18], we used a ResNet-50 [8] backbone pre-trained on Imagenet followed by a fully connected layer and $L_{2}$ normalization as our feature extractor. We trained for 15 epochs and perform early stopping based on average validation Prec@1. Other training hyperparameters are kept the same as CDS [18]. In our approach using synthetic data, we generated one synthetic example per real example in the training set. For full training details, please refer to Appendix C.
4.1 Cross Domain Retrieval Results
@@
Methods for Comparison. We compare 4 versions of SynCDR using different synthetic data generation methods. We also compare them to prior work using no synthetic data. CDS is the method proposed by Kim et al. [18] which uses a combination of in-domain instance discrimination [45] (ID) and cross-domain matching. As another comparison, we drop the latter of these two criteria and train only with in-domain ID loss. As we shall see, in the absence of same category examples across domains, this leads to better performance. Finally, we also report the performance of using simply an ImageNet-1K pre-trained ResNet-50 backbone, which serves as the initialization for all the other methods.
@@
We digress briefly to note that while large image embedding models based on vision transformers [5] and CLIP [25] pre-training are available [16] and provide robust image embeddings across domains, they may be expensive to store and run on smaller devices at inference time. For instance, a CLIP ViT-B model (the smallest CLIP available) has 160M parameters compared to a ResNet-50’s 25M. We thus only use large pre-trained models at training time. Although knowledge distillation [10] is possible from these large CLIP models to smaller ones, as shown by Sun et al. [36] (in Fig. 3 of their paper), it still requires a large number of images and text captions (a few tens of millions). In Sec. A.3, we show that distillation is not effective with the limited size of our training datasets (which also have no captions available).
@@
Each of Tabs. 1, LABEL:, 2, LABEL: and 3 reports the results of cross-domain retrieval (Prec@1) across all pairs of domains for a given dataset. Each cell (except the ones for the pre-trained models) reports mean and standard deviation over 3 different training runs. The “Average” column reports the mean and the pooled standard deviation over the other columns in the table.
@@
[…]
@@
CDS - 33.4 ± 0.7 45.9 ± 0.8 51.4 ± 0.8 38.9 ± 1.2 41.3 ± 0.7 63.3 ± 1.3 45.7 ± 1.0
In-domain ID - 36.0 ± 0.8 49.0 ± 0.8 54.9 ± 0.5 40.0 ± 1.0 43.5 ± 0.6 67.0 ± 0.6 48.4 ± 0.7
SynCDR (Ours) CUT 36.7 ± 1.0 48.8 ± 1.0 54.3 ± 1.4 41.1 ± 0.6 45.0 ± 0.8 67.5 ± 0.5 48.9 ± 0.9
Img2Img 36.3 ± 0.7 48.7 ± 0.8 54.5 ± 1.3 40.9 ± 0.8 43.5 ± 0.8 67.7 ± 0.4 48.6 ± 0.8
InstructPix2Pix 37.9 ± 0.9 49.6 ± 1.0 54.5 ± 1.1 42.8 ± 0.6 46.3 ± 1.1 67.5 ± 0.4 49.8 ± 0.9
ELITE 37.1 ± 1.0 49.1 ± 1.0 54.8 ± 1.1 43.6 ± 0.5 48.3 ± 0.4 67.5 ± 0.9 50.1 ± 0.9
Table 3: Cross Domain Retrieval Prec@1 in different scenarios of Office-Home. We find synthetic data helps significantly in cases with Clipart domain, but not so much in other cases, where Imagenet pre-training is already quite good. ELITE data is again best on average.
@@
We also find that CUT and InstructPix2Pix perform at par with each other on average, but each is better in different scenarios. InstructPix2Pix is poorer in scenarios involving sketch since it is worse at “filling in color” into pencil sketches (as seen from Fig. 2), when simply prompted to convert a sketch to a painting or a clipart. On the other hand, CUT can add color to sketches based on it’s training on the two domains, even though these colors may not always be correct. In the clipart-painting scenario, InstructPix2Pix has a bigger edge in the quality of data being able to rely on Stable Diffusion’s prior understanding of these domains (more examples in Appendix D). Finally, we see that synthetic data from ELITE performs the best. Aside from generating good quality data, it does not suffer from InstructPix2Pix’s limitations in adding color to sketches since ELITE parses the object into a textual token and then generates a painting of it, without having to be restricted by the amount of edit between the original and the generated image.
@@
Office-Home. The results for this dataset are in Tab. 3. We first note than neither CDS not in-domain ID could improve performance over ImageNet pre-training (this was also found about In-Domain ID in [18], in the scenario where both source and target have data from all categories). This is likely because of the relatively small size of the dataset and similarity of the objects to Imagenet. That said, synthetic data does help overall, with ELITE being best generation method. We observe that performance significantly improves in cases where Clipart is involved, possibly since it is significantly distinct from the other three visual domains. All of other domains have a significant number of realistic photographs (often times sharing very similar images), and hence an Imagenet pre-trained network serves as a good retrieval model for these domains.

### 19 Mar 2024 v2

@@ Abstract
In cross-domain retrieval, a model is required to identify images from the same semantic category across two visual domains. For instance, given a sketch of an object, a model needs to retrieve a real image of it from an online store’s catalog. A standard approach for such a problem is learning a feature space of images where Euclidean distances reflect similarity. Even without human annotations, which may be expensive to acquire, prior methods function reasonably well using unlabeled images for training. Our problem constraint takes this further to scenarios where the two domains do not necessarily share any common categories in training data. One instance where this can occur is when the two domains in question come from different versions of some biometric sensor recording identities of different people. We posit a simple solution, which is to generate synthetic data to fill in these missing category examples across domains. This, we do via category preserving translation of images from one visual domain to another. We compare approaches specifically trained for this translation for a pair of domains, as well as those that can use large-scale pre-trained text-to-image diffusion models via prompts, and find that the latter can generate better replacement synthetic data, leading to more accurate cross-domain. Our best SynCDR model can outperform prior art by up to 15%. Code for our work is available at https://github.com/samarth4149/SynCDR
Keywords:
Cross-domain Retrieval Synthetic Data Diffusion Models
@@ 1 Introduction
Figure 1: Motivation. Cross Domain Retrieval problems show up in many applications, and prior work has developed solutions in different scenarios, including when labeled data is absent [22]. What these solutions rely on however is the presence of same category data in both domains so similar pairs can be discovered. When these are missing, such approaches can fail. Our solution is to make up for these missing examples using synthetic data (generated via label preserving translation). While such data may not be a perfect replacement (e.g. the real image generated from the painting in the above example may not be entirely realistic because of a white background), we show that they are still useful for training cross-domain retrieval models.
@@ 2 Related Work
Self-supervised Feature Learning for Cross-Domain Retrieval. While approaches above rely on labels for training, these may be expensive in labor-hours to collect. In the absence of any labels representation learning can be done by optimizing for different self-supervision objectives [10, 4, 11, 5]. Kim et al. [22], followed by others [51, 15], demonstrated that category discriminative domain-adaptive features can be learned without the presence of any labels using self-supervised criteria within and across domains. This, however, relies on the presence of same categories in both domains in training data. If unavailable, the cross-domain self-supervision losses can degrade retrieval performance by training the model to match unrelated instances. This is the problem we tackle in this paper, by using synthetically generated examples to make up for the missing categories of data.
@@ 3 Approach
Notation. Let the two domains be A and B. At training time, two sets of images are available (with no accompanying labels), $\mathcal{D}_{A}=\{\boldsymbol{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{D}_{B}=\{\boldsymbol{x}^{B}_{i}\}_{i=1}^{n_{B}}$. The ground truth category labels are unknown but let each image in $\mathcal{D}_{A}$ come from a category label in $Y_{A}$, and each image in $\mathcal{D}_{B}$ come from a category label in $Y_{B}$. It is known that $Y_{A}\cap Y_{B}=\phi$ (This is not a strict requirement for SynCDR to function and we report results where $Y_{A}$ and $Y_{B}$ share some categories in Sec. 4.4). The outcome of training is a feature extractor $f:\mathcal{X}\rightarrow\mathbb{S}^{d}$ (where $\mathcal{X}$ is the support of all images, and $\mathbb{S}^{d}$ the $d$-dimensional unit hypersphere $\mathbb{S}^{d}=\{\boldsymbol{x};\boldsymbol{x}\in\mathbb{R}^{d},\left|\left|x\right|\right|_{2}=1\}$), such that cosine similarities in the output feature space reflect category semantics.
For evaluation are provided two sets $\mathcal{D}^{(test)}_{A}$ and $\mathcal{D}^{(test)}_{B}$ both of which have images with known category labels and each set has at least one image from each label in $Y_{A}\cup Y_{B}$. At test time, given an image $\boldsymbol{x}^{A}$ in $\mathcal{D}^{(test)}_{A}$, it is desired to fetch images of the same category label from $\mathcal{D}^{(test)}_{B}$, and vice-versa given an image in domain B. This is done by ranking in decreasing order $\{f(\boldsymbol{x}^{A})^{\top}f(\boldsymbol{x}^{B});\boldsymbol{x}^{B}\in\mathcal{D}^{(test)}_{B}\}$ and picking top K to calculate precision@K (for different values of K).
@@
$L_{PPP}(X_{A},X_{syn,B})$ $=\frac{1}{m}\sum_{\boldsymbol{x}^{A}\in X_{A}}-\log\left(\frac{\exp(f(\boldsymbol{x}^{A})^{\top}f(\boldsymbol{x}_{syn}^{B}))}{\sum_{\boldsymbol{x}\in X_{syn,B}}\exp(f(\boldsymbol{x}^{A})^{\top}f(\boldsymbol{x}))}\right)$
@@
Training SynCDR. For training we combine $L_{PPP}$ with the CDS criterion [22] ($L_{CDS}$), and minimize the following loss via minibatch stochastic gradient descent. To meet length constraints, we excluded a precise definition of $L_{CDS}$ and refer readers to Sec. 3 and Fig. 3 of [22]. For the sake of current discussion, we simply describe it as a combination of in-domain and cross-domain contrastive criteria given two sets of unlabeled images from two domains.
@@
Figure 2: Synthetic examples from different translation methods. We compare synthetic data generated using 4 methods. Since ELITE is not specifically restricted to closely mimicking the original image, we can generate more natural examples of target domain data, which serves as better synthetic replacement for missing real data. For more discussion, refer to Sec. 3.1.
@@ 3.1 Methods for Synthetic Data Generation
Img2Img/SDEdit [24] uses a pre-trained Stable Diffusion model for image editing by partially following the diffusion forward noising process (so that not all content in the input image is lost due to noising) and then denoising based on a text prompt. For instance in the example in Fig. 2, for translating real images of birds from the CUB dataset to paintings, we used the prompt “A painting of a bird.” and for translating sketches to paintings for Domainnet objects, we used “A painting of an object.”. The method allows for specification of edit strength $\in[0,1]$, which is the fraction of diffusion noising steps to follow. The value 00 leads to no edit and $1$ leads to completely following the prompt and retaining none of the input image information. In our experiments, we set this parameter to $0.3$ based on validation performance. For prompts used for all datasets and other generation parameters, we refer readers to Appendix 0.B.
@@
InstructPix2Pix [2] is a Stable-Diffusion model fine-tuned and adapted for following natural language instructions for editing images. For converting to paintings (for both photos of birds as well as sketches of objects from Domainnet), we input the original image and the prompt “Convert to a painting.”. With InstructPix2Pix, we can provide a parameter $image\_guidance\_scale\geq 1$. The larger the value of this parameter, more is the output image similar to the input image. In our experiments, we set this parameter to $1.2$ based on validation performance. Full prompts and other generation parameters for all datasets and additional examples are in Appendix 0.B.
@@
The above translation methods attempt to closely resemble the original image in translations and we find that this is often not enough to accurately represent data in the other domain. In such scenarios, images may need additional content to make it look like the target domain. For instance, in Fig. 2, while the above methods do a decent job of converting paintings to pencil sketches or real photos to paintings, the translation in the opposite direction is much poorer. In translating a sketch to a painting (row 2), Img2Img and InstructPix2Pix fail to add color and CUT adds some incorrect colors based on what it could learn from unpaired images. Similar is the case in converting a painting of a bird to a realistic photograph (row 4), where a background may be expected but above methods fail to generate one (CUT does generate one, but does not generate an natural looking photograph). While Img2Img and InstructPix2Pix allow for increasing the amount of edit via a controllable parameter, in Appendix 0.B, we show that this often leads to non-preservation of image category. The following generation method can avoid these issues.
@@
ELITE [47] is a personalization approach based on Stable Diffusion and a CLIP-based image encoder. The encoder encodes an object instance from an image as a word in Stable Diffusion’s vocabulary, which can then be used for generation with prompts. In the example in Fig. 2, for generating a painting (from either a real photo or a sketch), we input the image to ELITE’s image encoder and use the prompt “A painting of $\langle S\rangle$.” (where $\langle S\rangle$ is the image encoded as a word, or more precisely, a token embedding). For full details, refer to Appendix 0.B.
@@ 4 Experiments
Datasets and Evaluation. We ran experiments on three different datasets : DomainNet [26], CUB [45, 46] and Office-Home [44]. We experimented with the 126 class subset of DomainNet used in [22] with 3 domains : clipart, painting and sketch. For CUB, we use two domains : Real (coming from the CUB-200-2011 data [45]) and Paintings (coming from [46]). In CUB, we removed categories that had fewer than 5 paintings, leaving us with 187 categories. The Office-Home dataset has 65 categories in 4 domains : Art, Clipart, Product and Real World.
We experimented with each possible pair of domains in each dataset. For the experiments, we made a 50:20:30 split of the data in the two domains using the first for training, second for validation and third for testing. For each pair of domains we split the training data into two disjoint sets of categories (of equal size) and run two experiments : one where category set 1 is from domain A and category set 2 is from domain B, and the other where this is reversed. Note that testing in either of these cases takes place on the same data, and we report average Prec@1 of retrieval in either direction (given domain A image, fetch a domain B image of same category and vice-versa), and the overall average of these over all pairs of domains in the dataset. In Sec. 0.A.3, we also report Prec@5 and Prec@15.
Other Implementation Details. Similar to CDS [22], we used a ResNet-50 [12] backbone pre-trained on Imagenet followed by a fully connected layer and $L_{2}$ normalization as our feature extractor. We trained for 15 epochs and perform early stopping based on average validation Prec@1. Other training hyperparameters are kept the same as CDS [22]. In our approach using synthetic data, we generated one synthetic example per real example in the training set. For full training details, please refer to Appendix 0.C.
4.1 Cross Domain Retrieval Results
Table 1: Cross Domain Retrieval Prec@1 in different scenarios of DomainNet. As highlighted in Fig. 1, prior art (CDS) cannot perform well (and is only about as good as a pre-trained ImageNet model) in this scenario where training data does not have examples from the same categories across domains. The cross-domain criterion from CDS, impedes performance and using only the In-domain ID criterion for training performs better. As seen in the results, synthetic data can make up for missing real examples and lead to improved performance. ELITE provides the best quality synthetic data for SynCDR, resulting in best cross-domain retrieval performance.
@@ 4.1 Cross Domain Retrieval Results
Methods for Comparison. We compare 4 versions of SynCDR using different synthetic data generation methods. We also compare them to prior work using no synthetic data. CDS is the method proposed by Kim et al. [22] which uses a combination of in-domain instance discrimination [49] (ID) and cross-domain matching. As another comparison, we drop the latter of these two criteria and train only with in-domain ID loss. As we shall see, in the absence of same category examples across domains, this can lead to better performance. Finally, we also report the performance of using simply an ImageNet-1K pre-trained ResNet-50 backbone, which serves as the initialization for all the other methods.
@@
We digress briefly to note that while large image embedding models based on vision transformers [7] and CLIP [30] pre-training are available [20] and provide robust image embeddings across domains, they may be expensive to store and run on smaller devices at inference time. For instance, a CLIP ViT-B model (the smallest CLIP available) has 160M parameters compared to a ResNet-50’s 25M. We thus only use large pre-trained models at training time. Although knowledge distillation [14] is possible from these large CLIP models to smaller ones, as shown by Sun et al. [41] (in Fig. 3 of their paper), it still requires a large number of images and text captions (a few tens of millions). In Sec. 0.A.2, we show that distillation is not effective with the limited size of our training datasets (which also have no captions available).
@@
Each of Tabs. 1, LABEL:, 2, LABEL: and 3 reports the results of cross-domain retrieval (Prec@1) across all pairs of domains for a given dataset. Each cell (except the ones for the pre-trained models) reports mean and standard deviation over 3 different training runs. The “Average” column reports the mean and the pooled standard deviation over the other columns in the table. We additionally report Prec@5 and Prec@15 in Sec. 0.A.3, while noting that they follow the same trends as seen in Prec@1.
[…]
@@
Table 3: Cross Domain Retrieval Prec@1 in different scenarios of Office-Home. As with other benchmarks, we find that SynCDR with ELITE generated data results in best performance on average over all scenarios in Office-Home. We see synthetic data results in relatively small performance improvements in scenarios with small domain gaps (e.g. Product-Real), but more significant improvements in larger domain gap scenarios (e.g. those involving Clipart). Refer to Sec. 4.1 for further discussion.
@@
CDS - 36.5 ± 0.3 50.0 ± 0.4 54.9 ± 0.3 43.0 ± 0.2 45.2 ± 0.3 69.1 ± 0.2 49.8 ± 0.3
In-domain ID - 36.3 ± 0.3 50.0 ± 0.4 54.7 ± 0.3 43.4 ± 0.4 45.4 ± 0.4 69.0 ± 0.3 49.8 ± 0.3
SynCDR (Ours) CUT 38.3 ± 0.4 50.8 ± 0.2 55.5 ± 0.6 43.3 ± 0.5 46.6 ± 0.3 69.3 ± 0.3 50.6 ± 0.4
Img2Img 36.4 ± 0.3 49.9 ± 0.3 54.9 ± 0.5 43.4 ± 0.4 45.2 ± 0.4 69.0 ± 0.3 49.8 ± 0.4
InstructPix2Pix 38.6 ± 0.5 50.2 ± 0.2 54.7 ± 0.3 43.9 ± 0.2 46.7 ± 0.4 69.0 ± 0.2 50.5 ± 0.3
ELITE 38.3 ± 0.5 51.3 ± 0.3 55.4 ± 0.3 44.1 ± 0.6 46.3 ± 0.5 69.3 ± 0.3 50.8 ± 0.4
@@
We also find that CUT and InstructPix2Pix perform at par with each other on average, but each is better in different scenarios. InstructPix2Pix is poorer in scenarios involving sketch since it is worse at “filling in color” into pencil sketches (as seen from Fig. 2), when simply prompted to convert a sketch to a painting or a clipart. On the other hand, CUT can add color to sketches based on it’s training on the two domains, even though these colors may not always be correct. In the clipart-painting scenario, InstructPix2Pix has a bigger edge in the quality of data being able to rely on Stable Diffusion’s prior understanding of these domains (more examples in Appendix 0.E). Finally, we see that synthetic data from ELITE performs the best. Aside from generating good quality data, it does not suffer from InstructPix2Pix’s limitations in adding color to sketches since ELITE parses the object into a textual token and then generates a painting of it, without having to be restricted by the amount of edit between the original and the generated image.
@@
Office-Home. The results for this dataset are in Tab. 3. As with other benchmarks, we found that SynCDR with ELITE’s synthetic data resulted in best cross-domain retrieval performance overall. We find on some domains similar to ImageNet data (such as Product-Real), an ImageNet pre-trained backbone performs quite well, and CDS and in-domain ID cannot improve much on this given they are only self-supervised training objectives functioning with no dataset specific labels. Additionally, we find that on such small domain gap scenarios (Product has realistic images on a white background, and some training images were even found to occur in both Product and Real domains), performance improvement from synthetic data is relatively small. This is because the problem of missing examples as shown in Fig. 1 is not as big with small domain gaps. On the other hand, when the domain gap is larger (e.g. when Clipart is involved) SynCDR helps improve performance more significantly.